A set of eight built-in bookshelves is to be constructed in a room with a floor-to-ceiling clearance of 9 ft 6 in. If each shelf is 3/4 in. thick and equal spacing is desired between shelves, what is the space between the shelves (assume there is no shelf against the ceiling and no shelf on the floor) Thank you so much.
Answer by richwmiller(17219)

8*3/4=6 inches for the shelves which leaves 9 feet for the spacing.
9*12=108 inches if we have 8 shelves then we have 108/8=27/2 =13.5 inches
or assuming that the first shelf is the same space off of the floor then there is a space of 12 inches between each shelf